OpenStudy (anonymous):

Determine the slope of the line that contains the points E(5,4) & F(8,1). Select one: a. -1 b. -3/13 c. undefined d. 5/13

OpenStudy (texaschic101):

slope formula = (y2 - y1) / (x2 - x1) (5,4)....x1 = 5 and y1 = 4 (8,1)...x2 = 8 and y2 = 1 now we sub.. slope = (1 - 4) / (8 - 5) slope = -3/3 slope = -1